OpenStudy (anonymous):

Please help! Simplify the expression. Assume all variables are non negative. Click to view :)

OpenStudy (anonymous):

\[\sqrt[3]{(64x^4y^5)}\]

OpenStudy (anonymous):

8ax^2y^2 root of y

OpenStudy (anonymous):

Actually i believe it is \[4xy \sqrt[3]{xy^2} \]

OpenStudy (anonymous):

Since it is a cube root

OpenStudy (anonymous):

It is [8ax ^{2}y ^{2}\sqrt{y}\]

OpenStudy (anonymous):

Hey you know what? I know a reliable tutor site and it is free http://lhh.tutor.com/?ProgramGUID=9a1970d7-0dee-4ef2-ac3b-8a0a6ccc09ef go there, get a tutor and ask. I do think I might be wrong so go ask there :) Hope I helped!!

OpenStudy (anonymous):

Just exit out the sign in thingy

OpenStudy (anonymous):

@Melodyism , the only reason i think you are wrong is because your factoring would be right if it were a square root

OpenStudy (anonymous):

You might be right IDK. lol did I sound mad or something? Sorry if I did I am not.. I did not even learn this yet :P Just tried it.. haha

OpenStudy (anonymous):

@Samreina

OpenStudy (anonymous):

It's fine, you did not sound mad. Just making sure there is no confusion.